# SynLlama: Generating Synthesizable Molecules and Their Analogs with Large Language Models

### v3

@@ 3.1 Synthesis Planning for Unseen Molecules
| Testing-Set 1 | SynNet48 | 107 | - | 107 | 0.46 |
| ChemProjector49 | 301 | - | 301 | 0.79 | |
| SynLlama(RXN Set 1) | 631 | 125 | 669 | 0.93 | |
| Testing-Set 2 | SynLlama(RXN Set 2) | 538 | 114 | 582 | 0.91 |
| Enamine Data | SynNet | 110 | - | 110 | 0.57 |
@@
| Synformer50 | 660 | - | 660 | 0.91 | |
| SynLlama(RXN Set 1) | 527 | 100 | 568 | 0.87 | |
| SynLlama(RXN Set 2) | 691 | 232 | 741 | 0.92 | |
@@
| SynLlama(RXN Set 1) | 165 | 95 | 223 | 0.66 | |
| SynLlama(RXN Set 2) | 197 | 152 | 287 | 0.68 | |
Having demonstrated that SynLlama models can reliably predict reaction sequences and building blocks, we now use SynLlama to plan the synthesis of two groups of previously unseen molecules. The first group comprises two sets of 1000 testing molecules, generated similarly to the training data using testing building blocks and corresponding reaction templates. The second group includes two sets of 1000 molecules extracted from the publicly available ChEMBL database19 and the Enamine diversity set20.
Initially, we follow the standard reconstruction approach within the Enamine chemical space: SynLlama generates raw output predictions in a retrosynthetic manner, and this output then gets processed through the reconstruction algorithm to create target input molecules or analogs using building blocks exclusively from the Enamine library. This evaluation method aligns with that used by existing approaches such as SynNet48, ChemProjector49, and Synformer50, and we compare these methods on the number of reconstructions of the target molecule with Enamine BBs in Table 2. In this comparison, SynNet48 serves as a baseline for synthesizable chemical space coverage.49 Furthermore, when trained on their respective reaction sets and using just Enamine BBs along, SynLlama outperforms ChemProjector49 and matches the performance of Synformer50, while reducing the number of training data by 60-fold and 40-fold.
@@
In addition to reducing data requirements for training, SynLlama can also reconstruct target molecules using commercially available BBs beyond Enamine, even without specific training for this purpose. These ‘New BBs’ add possible synthetic pathways to reconstruct target molecules in all four datasets, as seen in Table 2 and detailed in Tables S3 and S4. Since some target molecules can be synthesized through multiple pathways, either using only Enamine BBs or with the addition of New BBs, the ‘Total’ column in the table reflects the number of unique target molecules reconstructed with SynLlama. With these New BBs, SynLlama’s best reconstruction rates increase to 66.9%, 58.2%, 74.1%, and 28.7% for the 1000 molecules in both Testing Sets, Enamine, and ChEMBL data, respectively. These results clearly outperform previous methods that lack such search capacity of novel building blocks. It is especially noteworthy that New BBs significantly enhance synthetic accessibility of unseen molecules, suggesting that SynLlama learns reaction chemistry well enough to predict novel building blocks.
@@
Finally, when the target molecule cannot be reconstructed, we assess the quality of the analog using three molecular similarity scores between the target molecule and its most similar analog: (1) Tanimoto similarity based on 4096-bit Morgan fingerprints64, (2) 4096-bit Morgan fingerprints of Murcko scaffolds65, and (3) Gobbi 2D pharmacophore fingerprints66. In table 2, the similarity metrics reported are average values of all trial molecules, including those that are fully reconstructed. In 2, we only show the first similarity metric due to the space limit. More details on the metrics and results for molecules that cannot be reconstructed are explored in Tables S5 and S6. Overall, these results collectively show that SynLlama is capable of planning synthesis for unseen molecules or their analogs with purchasable building blocks, highlighting its potential use in drug discovery.
@@ 3.2 Synthesizable Analog Search for De Novo Molecules
Previous research has shown that molecules proposed by generative models often face challenges in practical synthesizability39, 40, 41, even when optimized using empirical heuristics for synthetic accessibility. In this section, we demonstrate SynLlama’s potential to bridge the gap between generative molecule design and practical synthesis. Specifically, SynLlama is integrated with candidates from a structure-based generative model, iMiner30, which optimizes drug-likeness and molecular docking scores against the SARS-CoV-2 Main Protease (Mpro). More details in regards to iMiner are provided in the Supplementary Information.
Here, we select the top 500 iMiner-generated binders based on their AutoDock Vina67 docking scores to Mpro (PDB: 7L1168), and process them through SynLlama, trained on RXN Set 2, to generate synthesizable analogs constrained to Enamine BBs. For the top-10 generated analogs, we perform molecular docking with the same procedure and report their average docking score and average Morgan fingerprint similarity in Figure2. We choose the SynLlama model trained on RXN Set 2 for their superior performance in analog generation explored in the previous section. A similar distribution analysis of the iMiner analogs generated with SynLlama trained on RXN Set 1 is presented in Figure S3.
In Figure 2(a), the RMSE of docking scores of the generated analogs with respect to the original iMiner molecules is 1.04 kcal/mol, which is an acceptable range of inherent docking score errors reported by Trott et al67. Furthermore, as demonstrated in Figure 2(b), the SA score distribution of the generated analogs show a notable decrease from de novo generated binders, demonstrating improved synthetic accessibility of the analogs proposed by SynLlama. These results suggest that SynLlama can generate synthetic pathways for iMiner-proposed molecules or their analogs, enhancing synthetic accessibility of de novo molecules when combined with generative models.
@@ 3.3 Local Hit Expansion for Binder Molecules
Because SynLlama breaks down the original target molecule for synthesis into building blocks, by nature this method allows diverse exploration around parts of the molecular scaffold rather than only on a whole target molecule. In a final task, we apply SynLlama to expand on a hit molecule for the same SARS-CoV-2 Mpro protein (PDB: 7LTJ)69 to discover synthesizable molecules that have better relative binding free energies (RBFEs) confirmed by both experiments and accurate free energy perturbation (FEP) calculations.
@@
As shown in Figure 3(a), the hit molecule has a core scaffold of uracil and ortho-dichlorobenzene connected by a piperizine linker. Inspired by an experimental hit expansion campaign by Kneller et al. 70, we follow their practices to propose only functional group substitutions on the benzene ring while keeping the linker and uracil intact. Specifically, we use SynLlama to generate 50 synthesizable analogs constrained to only Enamine BBs of the hit compound and filter for molecules that only modifies the benzene ring, harvesting a total of 8 analog molecules that fulfill the criteria. The analogs are placed in a pose configuration similar to the original hit molecule and we perform FEP on all 8 molecules.
To verify the FEP results, we first choose around 10 synthesized and experimentally tested molecules to benchmark the accuracy of FEP for this system. In Figure S4, all calculated FEP values show a good correlation with experimental converted from IC50, with an average RMSE of less than . After this validation, we run FEP for the eight proposed molecules to assess their binding affinities. As shown Figure 3(b), SynLlama rediscovers two molecules from the reported hit expansion campaign, both with experimentally validated IC50 values. Notably, one molecule highlighted in blue was identified as the best compound from the expansion efforts reported by Kneller et al.70. The other six molecules highlighted in green also demonstrate desirable FEP-calculated binding affinities compared to the lead molecule, suggesting new directions for hit optimization in this ligand series. These results show that SynLlama effectively explores local chemistry with purchasable building blocks, saving time and effort for medicinal chemists who would otherwise manually enumerate and check availability at different steps.
4 Discussion
@@
Motivated by recent advances in LLMs for chemistry54, 52, 38, 53, we aim to leverage data-efficient supervised fine-tuning (SFT) to transform the general-purpose Meta Llama 3 into SynLlama, an LLM-based generator capable of proposing synthesizable molecules and deducing synthetic routes for target molecules or their close analogs. Throughout the study, we successfully show that SynLlama can effectively explore a custom-defined chemical search space composed of around 230,000 Enamine building blocks (BBs) and well-validated organic reactions (RXNs), after it has been fine-tuned on synthetic pathway data sampled from this specified chemical space. What’s more, despite utilizing nearly two orders of magnitude fewer synthetic pathways, SynLlama exhibits strong performance in key drug discovery tasks compared to existing models. Specifically, we have demonstrated that SynLlama can effectively aid in various stages of drug discovery that include synthesis planning, synthesizable analog generation for de novo molecules, and local hit expansion.
@@
Because SynLlama is built on a general-purpose LLM instead of training from scratch48, 49, 50, it offers a number of unique advantages and possibilities for further improvement. For example, when generating our fine-tuning data, we sampled from the predefined chemical search space of 230K Enamine building blocks (BBs) and two sets of reaction templates (RXNs), but we did not embed these extensive requirements in the context window of the LLM. As a result, for our largest dataset with 2 million synthetic pathways, the model only saw each BB a few dozen times, while each RXN template appeared hundreds of thousands of times. Consequently, while SynLlama efficiently memorizes the allowed RXNs, it only captures the distribution of Enamine BBs, which enables SynLlama to extrapolate to unseen yet purchasable building blocks outside of Enamine. This generative ability surpasses other existing methods, which typically formulate the synthetic pathway generation as a Markov decision process within a defined chemical space.
@@
In addition to its ability to extrapolate outside the training chemical space, the underlying Llama-3.2-1B used by SynLlama is relatively small and more predictive power would be expected if we train on larger LLMs with more data and compute power. However, we observe that a smaller LLM with fewer parameters can be turned into an expert model for complex tasks after SFT with sufficient data. This opens up opportunities to employ smaller expert models for various chemical tasks, benefiting from faster inference speeds, which can make these models more desirable. Moreover, optimal hyperparameters like temperature and top-p can vary between training and inference phases, depending on the downstream tasks. During inference, most valid raw outputs are generated under relatively low temperature and top-p settings. However, when the model is paired with reconstruction algorithms that require less strict adherence to reaction chemistry, higher temperature and top-p values can be used. This allows for a broader exploration of the Enamine chemical space, enabling the generation of more diverse and relevant analogs. This property is especially desirable in tasks that require extensive exploration, such as the hit expansion example we demonstrate. Another exciting success is that when SynLlama is paired with another generative model, the generated analogs can maintain docking scores while exhibiting a shift in synthetic accessibility scores. This result suggests that SynLlama can serve effectively as a post-processor for other de novo generative models, ensuring the production of more synthesizable compounds with clear reaction pathways.

### v4

@@ 3 Results
We examine SynLlama’s performance in synthesis planning of a diverse set of previously unseen compounds. We also explore the utility of the SynLlama workflow in real-world drug discovery applications, including its integration with generative algorithms to enhance the synthetic accessibility of proposed molecules while preserving their chemical properties and expanding the library of active compounds in the defined synthesizable space with as good or improved binding affinity metrics.
@@ 3.1 Synthesis Planning for Unseen Molecules
Having demonstrated that SynLlama models can reliably predict reaction sequences and building blocks in Table 1, we now use SynLlama to plan the synthesis of two groups of 1000 previously unseen molecules from the Enamine Diversity Set9 and the publicly available ChEMBL database8. These datasets are specific to drug-like molecules, unlike the training data used for SynLlama; the drug-related property distribution of both sets of molecules against the training data is shown in Figure S3. In this validation, we test whether known synthesizable molecules can be reconstructed accurately from the baseline and SynLlama models as summarized in Table 2.
We first consider the standard reconstruction approach used by algorithms such as SynNet54, ChemProjector55, and Synformer56 to create target molecules or analogs using BBs exclusively from the Enamine library. In this comparison, SynNet54 and ChemProjector55 serve as a baseline comparison for synthesizable chemical space coverage for RXN 1, whereas Synformer56 is the baseline comparison on the expanded RXN 2 templates. As seen in the first column of Table 2 and Supplementary Table S3, when trained with their respective reaction sets and using only Enamine BBs, SynLlama outperforms all three methods while reducing the number of training data by 40-to 60-fold.
Although SynLlama yields higher percentages of successful ChEMBL reconstructions compared to SynNet and ChemProjector, and is on par with Synformer when only using Enamine BBs, there is a degradation of performance across all methods for ChEMBL compared to the Enamine Diversity set reconstructions. We attempted to improve upon the ChEMBL result by reformulating the training reaction data with extra filtering such that the product molecule distributions conform to a similar drug-like property distribution as the ChEMBL set as seen in Supplementary Figure S4. We then performed supervised fine-tuning following the same procedure as described in Section 2, but now with this filtered set of training data, in hope of better generating synthetic pathways for molecules with drug-like properties. However, as shown in Supplementary Table S3, there is now a performance loss over both the Enamine Diversity and ChEMBL drug-like targets. This result suggests that training on more diverse product molecules ultimately benefits synthesis planning for the drug-like targets more than specializing the LLM further with more restrictive training data. In addition, the curated ChEMBL data appears unique and outside the synthesizable chemical space made up of Enamine BBs and RXN templates.
@@
| Enamine Diversity Set | SynNet | 110 | - | 110 | 0.57 |
@@
| Synformer | 660 | - | 660 | 0.91 | |
| SynLlama(RXN 1) | 527 | 100 | 568 | 0.87 | |
| SynLlama(RXN 2) | 691 | 232 | 741 | 0.92 | |
@@
| SynLlama(RXN 1) | 165 | 95 | 223 | 0.66 | |
| SynLlama(RXN 2) | 197 | 152 | 287 | 0.68 | |
@@
However, unlike baseline methods that only generate molecules using Enamine BBs, SynLlama has the extra capacity of reconstruct target molecules with commercially available BBs beyond Enamine due to its generative capabilities, even without specific training for this purpose. As seen in Table 2 the ‘New BBs’, restricted to those purchasable through Molport69, add possible synthetic pathways to reconstruct target molecules in all datasets and RXN templates. This also helps the ChEMBL data as well given that the molecules generated with the new BBs remain drug-like (Supplementary Figure S5). Since some target molecules can be synthesized through multiple pathways, either using only Enamine BBs or with the addition of New BBs, the ‘Total’ column in Table 2 reflects the number of unique target molecules reconstructed with SynLlama. With these New BBs, SynLlama’s best reconstruction rates increase to 74.1% for the 1000 molecules in the Enamine Diversity Set, and encouragingly to 28.7% for the ChEMBL data. These results show that SynLlama learns reaction chemistry such that it can predict novel BBs to increase synthetic accessibility.
@@
When the target molecule cannot be reconstructed, we assess the quality of the analog using a molecular similarity score between the target molecule and its most similar analog using Tanimoto similarity based on 4096-bit Morgan fingerprints70. In Table 2, the similarity metrics reported are average values of all generated molecules, including target molecules that are fully reconstructed (with a score of 1). Tables S6 and S7 also provide similarity metrics based on the 4096-bit Morgan fingerprints of Murcko scaffolds71, and Gobbi 2D pharmacophore fingerprints72, while including or excluding target molecules that are fully reconstructed. Overall, these results collectively show that SynLlama is highly capable of planning synthesis for related analog molecules with very good similarity, aided most by increased synthetic pathways using purchasable building blocks.
Finally, Supplementary Table S3 also considers reconstruction performances for the Enamine Diversity Set and ChEMBL Data based on forward synthesis as opposed to retrosynthesis, and for test molecules derived from tree-like synthesis pathways. It is evident that SynLlama performs better for retrosynthesis relative to forward synthesis that is used more successfully by Synformer, and retrosynthesis also performs better than the baseline methods for branching synthetic pathways.
@@ 3.2 Synthesizable Analog Search for De Novo Molecules
Previous research has shown that molecules proposed by generative models for binding to protein targets often face challenges in both reliability and practical synthesizability28, 29, 30. In particular, medicinal chemists are often reluctant to devote time and expensive resources to specialized synthesis of hit molecules due to the high false positive rates arising from the unreliability of docking scores in drug discovery. Instead finding closely related compounds that are constrained to Enamine BBs allows for an inexpensive purchase to verify hits before further refinements are deployed to gain lead drug molecules. In this section, we demonstrate SynLlama’s potential to bridge the gap between generative molecule design and practical synthesis for molecules that are optimized for drug-like applications such as binding assays.
Specifically, we consider two different generative methods for de novo drug molecules in Figure 2: a 1D-to-3D LSTM model, iMiner19, which optimizes drug-likeness and AutoDock-vina73 docking scores, and Pocket2Mol22, a 3D graph transformer model which also optimizes AutoDock-vina73 docking scores. Using iMiner we generated 500 molecules for the SARS-CoV-2 Main Protease (SARS2 MPro)74, and used Pocket2Mol to generate 500 molecules each for the protein targets Thrombin75 and TYK276, 77. The first interesting observation is that only 1% of generated molecules from both iMiner and Pocket2Mol can be synthetically reconstructed using SynLlama or ChemProjector, emphasizing that the generative models create difficult synthesis targets. Hence all 500 compounds for each protein target were then processed through SynLlama trained on RXN 2 to generate synthesizable analogs constrained to Enamine BBs. For all generated analogs we performed molecular docking with AutoDock-vina via using the same protocol as the original binders.
Figure 2(a,b) shows the RMSE of the docking scores between the target compounds and the generated analogs are 1.44 kcal/mol and 1.11 kcal/mol for Pocket2Mol and iMiner, respectively, both of which are within the acceptable range of inherent docking score errors reported by Trott et al73. Furthermore, as demonstrated in Figures 2(c,d), the SA score distribution of the SynLlama analogs generated for thrombin and TYK2 from Pocket2Mol show a notable improvement in synthetic accessibility at the expense of reduced similarity below the benchmarks in Table 2. SynLlama slightly improves SA for iMiner generated compounds for SARS2 MPro while maintaining a good similarity score on par with the benchmarks in Table 2. The fact that iMiner uses a drug-likeness score as part of its loss function19 may explain its better performance, or perhaps SARS2 MPro is an easier protein target than Pocket2Mol’s thrombin and TYK2 protein targets.
The better synthesizable analogs have good retention of the binding mode of the original generated molecules, visually confirmed by the representative docking poses for target-analog molecular pairs shown in Figure 3(a) for the three proteins. Figure 3(b,c) provides a few examples of the synthesis pathways for the analogs of the molecules derived from the generative molecules.
The final comparison consists of molecules that were identified as unsynthesizable by Gao et al30 via ASKCOS48. As also seen in Figure 2, the similarity scores of SynLlama analogs for this set are better than that observed for the Pocket2Mol generative model, and the generated analogs show a significant decrease in SA score, representing SynLlamas effective strategy of improving synthetic accessibility via analog generation. In Supplementary Figure S8, we also highlight the few concrete examples of target-analog pairs where objective scores are maintained while target synthetic accessibility scores are substantially reduced. Overall, these collective results highlight SynLlama’s utility in effectively proposing synthesizable analogs for de novo molecules, thereby enhancing their synthetic accessibility without compromising their desired drug-related properties.
@@ 3.3 Local Hit Expansion for Binder Molecules
Because SynLlama breaks down the original target molecule for synthesis into building blocks, by nature this method allows diverse exploration around parts of the molecular
@@
scaffold rather than only on a whole target molecule. In a final task, we apply SynLlama to expand on hit molecules for three protein targets used in the previous section, SARS2 Mpro78, Thrombin75, and TYK276, 77 to discover synthesizable molecules that have better relative binding free energies (RBFEs) confirmed by both experiments and accurate free energy perturbation (FEP) calculations.
@@
As shown in Figure 4(a), the hit molecule for SARS2 Mpro (7LTJ_Lead) has a core scaffold of uracil and ortho-dichlorobenzene connected by a piperizine linker. Inspired by an experimental hit expansion campaign by Kneller et al. 79, we follow their practice to propose only functional group substitutions on the benzene ring while keeping the linker and uracil intact. In Figures 4(c) and 4(e), we use the best-performing molecules from the Schrodinger FEP benchmarking set80 as our hits for Thrombin and TYK2. To expand on these hit molecules, we first identify the maximum common scaffolds among each group of molecules in the FEP benchmarking set and use the identified scaffolds to guide our selection of analog molecules. Specifically, we use SynLlama to generate 50 synthesizable analogs constrained to only Enamine BBs of the hit compound and filter for molecules that retain the scaffold. In the end, we harvest a total of 8, 14, and 11 analog molecules that fulfill the criteria for SARS2 Mpro, Thrombin, and TYK2, respectively. The analogs are then placed in a pose configuration similar to the original hit molecule for downstream FEP calculations.
To verify the FEP results, we first choose 10 synthesized and experimentally tested molecules to benchmark the accuracy of FEP for all three systems. In Supplementary Figure S9, all the calculated FEP values show a good correlation with the experimental converted from IC50, with an average RMSE of less than . After this validation, we run FEP for the all proposed molecules to assess their binding affinities. As shown in Figures 4(b), 4(d), and 4(f), a significant portion of the proposed analogs (7 of 8, 5 of 14, and 6 of 11, respectively) showed potency compared to their parent hits for SARS2 Mpro, Thrombin, and TYK2. These results successfully demonstrate that SynLlama can propose diverse yet potent analogs when constrained on a molecular scaffold.
Moreover, the fact that all suggested analogs also come with predicted pathways using common reaction templates and purchasable BBs from Enamine suggests SynLlama’s practical use for drug discovery. Because of this composite capacity of optimizing both potency and synthetic accessibility for small molecules, SynLlama successfully rediscovered Analog_1, the most potent compound reported by Kneller et al. in their hit expansion campaign for SARS2 Mpro79. Furthermore, for the Thrombin case, SynLlama explores the R2 substitution site, a region previously unaddressed by earlier molecular series, and demonstrates the generation of more potent molecules via FEP. These results confirm that SynLlama effectively explores local chemistry with readily available building blocks, providing a direct and efficient path for medicinal chemists to accelerate hit expansion.
4 Discussion and Conclusions
@@
Motivated by recent advances in LLMs for chemistry60, 58, 27, 59, we aim to leverage data-efficient supervised fine-tuning (SFT) to transform the general-purpose Meta Llama 3 into SynLlama, an LLM-based generator capable of proposing synthesizable molecules and deducing synthetic routes for target molecules or their close analogs. Throughout the study, we successfully show that SynLlama can effectively explore a custom-defined chemical search space composed of around 230,000 Enamine building blocks (BBs) and well-validated organic reactions (RXNs), after it has been fine-tuned on synthetic pathway data sampled from this specified chemical space. What’s more, despite utilizing nearly two orders of magnitude fewer synthetic pathways in training, SynLlama exhibits strong performance in key drug discovery tasks compared to existing models. Specifically, we have demonstrated that SynLlama can effectively aid in various stages of drug discovery that include synthesis planning, synthesizable analog generation for de novo molecules, and local hit expansion.
@@
Because SynLlama is built on a general-purpose LLM instead of training from scratch54, 55, 56, it offers a number of unique advantages and possibilities for further improvement. For example, when generating our fine-tuning data, we sampled from the predefined chemical search space of 230K Enamine building blocks (BBs) and two sets of reaction templates (RXNs), but we did not embed these extensive requirements in the context window of the LLM. As a result, for our largest dataset with only 2 million synthetic pathways, the model only saw each BB a few dozen times, while each RXN template appeared hundreds of thousands of times. Consequently, while SynLlama efficiently memorizes the allowed RXNs, it only captures the distribution of Enamine BBs, which enables SynLlama to extrapolate to unseen yet purchasable building blocks outside of Enamine. This generative ability surpasses other existing methods such as ChemProjector and Synformer that can only explore a predefined building block search space, such as within the Enamine Diversity Set9, and limits their ability to propose alternative synthesis pathway with novel building blocks.
@@
In addition to its ability to extrapolate outside the training chemical space, the underlying Llama-3.2-1B used by SynLlama is relatively small and more predictive power would be expected if we train on larger LLMs with more data and compute power. However, we observe that a smaller LLM with fewer parameters can be turned into an expert model for complex tasks after SFT with relatively little data. This opens up opportunities to employ smaller expert models for various chemical tasks, benefiting from faster inference speeds, which can make these models more desirable. Moreover, optimal hyperparameters like temperature and top-p can vary between training and inference phases, depending on the downstream tasks. During inference, most valid raw outputs are generated under relatively low temperature and top-p settings. However, when the model is paired with reconstruction algorithms that require less strict adherence to reaction chemistry, higher temperature and top-p values can be used. This allows for a broader exploration of the Enamine chemical space, enabling the generation of more diverse and relevant analogs. This property is especially desirable in tasks that require extensive exploration, such as the hit expansion example we demonstrate. Another exciting direction is coupling SynLlama with another generative model, in which we have shown generates analogs while maintaining good docking scores and simultaneously shifting to better synthetic accessibility scores. This result suggests that SynLlama can serve effectively as a post-processor for other de novo generative models, ensuring the production of more synthesizable compounds with clear reaction pathways.